OpenStudy (anonymous):

Find the inverse Laplace transform of F(s)=2/(s^2+3s-4).

zepdrix (zepdrix):

So looks like we need to factor the denominator, and then apply Partial Fraction Decomposition. Are you stuck on any particular part? :)

OpenStudy (anonymous):

After factoring the denominator, I don't know what to do. Can you help me step by step?

zepdrix (zepdrix):

Partial Fraction Decomposition tells us that the factors in the denominator can be broken down like this ~ Where A and B are unknown costants. \[\large \frac{2}{(s-1)(s+4)} \qquad=\qquad \frac{A}{s-1}+\frac{B}{s+4}\] From here, we'll multiply both sides by the denominator on the left giving us,\[\large 2=A(s+4)+B(s-1)\] Confused by anything yet?

OpenStudy (anonymous):

What's next?

zepdrix (zepdrix):

We'll plug in specific values for \(\large s\) in order to solve for A and B. We want values that will zero one of the unknowns. \[\large s=1 \qquad\to\qquad 2=A(1+4)+B(1-1) \qquad\to\qquad 2=5A+0\]So that get's us our A value. To get B we use the other convenient value for s. \[\large s=-4 \qquad\to\qquad 2=A(-4+4)+B(-4-1) \qquad\to\qquad 2=-5B\]

OpenStudy (anonymous):

Wait a minute, please don't leave.

zepdrix (zepdrix):

k :)

OpenStudy (anonymous):

And then?

zepdrix (zepdrix):

So we've found our A and B. \[\large A=\frac25 \qquad\qquad \qquad B=-\frac25\] Using our initial set up,\[\large \frac{2}{(s-1)(s+4)} \qquad=\qquad \frac{A}{s-1}+\frac{B}{s+4}\] That tells us that our fraction can be written like this,\[\large \frac{2}{(s-1)(s+4)} \qquad=\qquad \frac{\frac25}{s-1}-\frac{\frac25}{s+4}\]

zepdrix (zepdrix):

Let's pull those ugly fractions out of the numerator so it's a little easier to read.\[\large \frac{2}{5}\left(\frac{1}{s-1}\right)-\frac{2}{5}\left(\frac{1}{s+4}\right)\]

zepdrix (zepdrix):

These will be nice and easy to take the Inverse Laplace of now. Just need to remember back to your rules.\[\large \mathscr{L}\left[e^{at}\right] \qquad=\qquad \frac{1}{s-a}\] So let's apply this to our first term but in reverse,\[\large \mathscr{L}^{-1}\left[\frac{2}{5}\frac{1}{s-1}\right] \qquad=\qquad \frac{2}{5}\mathscr{L}^{-1}\left[\frac{1}{s-1}\right]\]

zepdrix (zepdrix):

Understand how to take the inverse of that first term there?

OpenStudy (anonymous):

Thank you so much for the help.

zepdrix (zepdrix):

No prob c: Partial Fractions can be a bit tough if you don't remember them. You might wanna do a few more of these to brush up on them.